ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 0, 8, 51 Chiu, Sou-yung, Chang-Hua University of Education, Taiwan Pedagogical Mathematics Knowledge, PMK and P(PMK)K Pedagogical Content Knowledge is a notion in teacher education, issued by L. S. Shulman. The content area concerned here is in mathematics and mathematics teaching, up to two levels, to be termed as PMK and P(PMK)K, respectively. The state of art of pedagogical mathematics knowledge is investigated via classroom observations, analysis of lesson plans, and problem-solving among math teachers in junior secondary schools and student teachers in their preparation years. Diagnosis of the lack of mathematical knowledge, including the understanding of concepts in school mathematics and deduction of mathematical propositions among teachers, helps teacher educators insight of the nature and the process of the task of incorporating pedagogical knowledge into the training program, wherein P(PMK)K is substantiated.